# Stepwise Alignment for Constrained Language Model Policy Optimization

### v1

@@ 7 Experiments
We empirically evaluate the effectiveness of SACPO in enhancing multiple criteria stepwise. This experiment focuses on improving the two criteria: helpfulness and safety (i.e., harmlessness), which aims to answer the following four questions:
  1. 1.
    Can SACPO enhance the helpfulness and safety of language model policies stepwise?
  2. 2.
    Can SACPO use different algorithms and datasets for each alignment?
  3. 3.
    How does $\beta/\lambda$ affect the balance between the usefulness and safety of the aligned model?
  4. 4.
    Can practical SACPO balance the usefulness and safety via model merging?
Experimental settings.
@@
We use the same experimental setup as in Safe RLHF wherever possible for fair comparisons. We employ the Alpaca-7B model as the SFT model. This model is trained to function as a conversational assistant with proficient ability in crafting helpful responses, which can generate both benign and harmful responses, making it an ideal candidate for allowing the effectiveness of alignment methods. We also utilize the PKU-SafeRLHF preference dataset [19], which comprises over 30,000 expert evaluations. Each record in this dataset presents a pair of responses to a specific prompt. These response pairs are ranked according to two distinct criteria: helpfulness and harmlessness. The harmlessness of a response is determined by its neutrality concerning 14 different risk categories. Conversely, the helpfulness of a response is judged based on factors such as its clarity, relevance, and overall quality.
Implementations. In this experiment, we apply DPO and KTO for each alignment on helpfulness and safety (i.e., harmlessness). Specifically, we implement the following four variants of SACPO:
  • DPO (helpfulness) $\rightarrow$ DPO (safety)
  • DPO (helpfulness) $\rightarrow$ KTO (safety)
  • KTO (helpfulness) $\rightarrow$ DPO (safety)
  • DPO (safety) $\rightarrow$ DPO (helpfulness)
@@
As for the parameter for KL divergence, we first set $\beta$ and then test a wide range of values of $\beta/\lambda$. As a result, we set $\beta=0.1$ when helpfulness is the first metric and $\beta=0.01$ otherwise. In addition, to evaluate the performance of the practical algorithm in Section 6, we implement linear model merging [42] between the helpfulness-aligned model and conservatively safety-realigned model with $\beta/\lambda=0.01$ trained by DPO (helpfulness) $\rightarrow$ DPO (safety). We use MergeKit [18] and tested three different mixing ratios; that is, $q=\{0.25,0.5,0.75\}$. For more implementation details (e.g., hyperparameters), see Appendix G.
@@
Evaluation. We use GPT-4 to measure the helpfulness and harmlessness (i.e., safety) of the responses generated by the LM policies. We base our prompts on those used in the Safe RLHF study with a slight adjustment in output format requirements to get more reliable evaluations (for more details, see Appendix G.2). Unlike the previous study, we employ two non-overlap sets of prompts to assess helpfulness and safety separately. We use prompts in the Safe RLHF study to evaluate safety, which has a high risk of inducing unsafe responses. For assessing helpfulness, we use prompts from the Alpaca-eval dataset ([24]) that are unlikely to result in harmful content. Using two sets of prompts, we aim to ensure that our assessments of helpfulness and safety did not affect each other.
(a) Win rate vs SFT model
(b) Elo score
[…]
Results. In Figure 2, we show the win rates of the LLMs trained by DPO (helpfulness) $\rightarrow$ DPO (safety) compared to the baseline SFT model (Figure 2) and Elo scores (Figure 2). The Elo scores are adjusted so that the baseline SFT model has a score of 1000. First, the model aligned for helpfulness in the first stage exhibited a significant improvement in helpfulness, as indicated by an increase in both the Elo score and win rate. However, this enhancement in helpfulness coincides with a reduction in safety. Subsequent alignment for safety in the second stage regains a substantial degree of safety with only a slight decrease in helpfulness. When compared with the model trained via Safe RLHF (i.e., beaver-7b-v1.0), DPO (helpfulness) $\rightarrow$ DPO (safety) yielded models (with $\beta/\lambda$ values of $\{0.01,0.025,0.05\}$) that outperforms in terms of both helpfulness and safety. Second, we observe that varying the $\beta/\lambda$ ratio allows us to fine-tune the equilibrium between helpfulness and safety. Specifically, by modifying $\beta/\lambda$, we could observe a near Pareto-front, demonstrating the trade-off between these two criteria. This result indicates the flexibility of the proposed method in obtaining a model with a desired trade-off between multiple criteria. However, DPO (helpfulness) $\rightarrow$ KTO (safety) performs significantly worse than DPO (helpfulness) $\rightarrow$ DPO (safety). We guess this is because KTO is inappropriate for the safety alignment, we will leave it to future work to identify the detailed reasons.
Figure 3 also shows that KTO (helpfulness) $\rightarrow$ DPO (safety) performs better than Safe RLHF, which demonstrates that different algorithms can be used for each alignment. This is a favorable result since only unpaired datasets are available in many scenarios. Finally, Figure 4 shows the experimental results of DPO (safety) $\rightarrow$ DPO (helpfulness) where alignments are executed in the order from safety to helpfulness. Its performance is comparable with Safe RLHF, but worse than DPO (helpfulness) $\rightarrow$ DPO (safety). This result implies that commutative law holds only partially and there is still a gap between theory and practice.
(a) Win rate vs SFT model
(b) Elo score
Figure 3: KTO (helpfulness) $\rightarrow$ DPO (safety). (a) Win rate against the SFT model. (b) Elo scores in harmlessness and helpfulness. While blue circles represent baselines, red circles represent SACPO where $\beta/\lambda$ is defined as a number in brackets.
(a) Win rate vs SFT model
(b) Elo score
@@
Figure 4: DPO (safety) $\rightarrow$ DPO (helpfulness). (a) Win rate against the SFT model. (b) Elo scores in harmlessness and helpfulness. Crosses represent baselines, and blue triangles represent DPO incorporating only safety. Red circles represent SACPO where $\beta/\lambda$ is defined as a number in brackets.
Finally, Figure 5 shows the effectiveness of our practical implementation of SACPO that is discussed in Section 6. Specifically, linear model merging allows us to balance helpfulness and harmlessness by averaging reward-aligned and conservatively safety-realigned policies without optimizing new ones, which contributes to reduced computational time and stable learning.
(a) Win rate vs SFT model
(b) Elo score
Figure 5: Practical implementation of SACPO. Green circles are LLMs that are linearly merged between DPO (helpfulness) and SACPO (DPO, 0.01).
@@ 8 Conclusion
We have introduced SACPO, a simple algorithm for constrained language model policy optimization. SACPO takes a stepwise approach that sequentially aligns an LM policy using off-the-self alignment algorithms (e.g., DPO, KTO) and datasets for each metric. This procedure is theoretically justified and provides us with many practical benefits such as simplicity, stability, and flexibility. Our theoretical results include the upper bounds regarding near-optimality and safety constraint violations, which bring insights into how to collect datasets. Empirically, SACPO performs better than Safe RLHF in enhancing helpfulness and harmlessness, and we further show the effectiveness of a practical variant based on linear model merging. We believe that SACPO reduces the barrier to aligning future LLMs to enhance the benefits of AI while minimizing any negative impacts.
@@ B.2 Lagrangian
B.3 Safety-aligned Model Can be Directly Obtained from Reward-aligned Model
@@
Appendix C Proof of Theorem 1
@@
$=$ $\ \frac{Z_{r^{\star}}(x;\pi_{\mathrm{ref}})}{Z_{r^{\star}+\lambda^{\star}g^{\star}}(x;\pi_{\mathrm{ref}})}\underbrace{\frac{1}{Z_{r^{\star}}(x;\pi_{\mathrm{ref}})}\pi_{\mathrm{ref}}(y\mid x)\exp\left(\frac{1}{\beta}r^{\star}(x,y)\right)}_{=\pi^{\star}_{r^{\star}}(y\mid x)}\exp\left(\frac{\lambda^{\star}}{\beta}g(x,y)\right)$
$=$ $\ \frac{1}{Y(x)}\pi^{\star}_{r^{\star}}(y\mid x)\exp\left(\frac{\lambda^{\star}}{\beta}g^{\star}(x,y)\right)$
@@
Appendix D Proof of Lemma 1 and Lemma 2
Proof.
Proof.
@@ Appendix E Proof of Lemma 3
Recall Assumption 2. By Lemma G.3 in Xiong et al. [43], the following inequality holds with probability at least $1-\delta$:
@@ F.1 Preliminary Lemmas
$\leq$ $\ c\cdot|\,g^{\star}(x,y)-\widehat{g}(x,y)\,|+|\,c-\widehat{\lambda}\,|\cdot|\,\widehat{g}(x,y)\,|$
@@ Appendix G Details of the experiments
G.1 Hyper-parameters
@@
DPO (helpfulness) $\rightarrow$ DPO (safety) $0.1$$\{0.01,0.025,0.05,0.1\}$
DPO (helpfulness) $\rightarrow$ KTO (safety) $0.1$$\{0.01,0.025,0.05,0.1\}$
KTO (helpfulness) $\rightarrow$ DPO (safety) $0.1$$\{0.01,0.025,0.05,0.1\}$
DPO (safety) $\rightarrow$ DPO (helpfulness) $0.01$$\{0.01,0.025,0.05,0.1\}$
G.2 GPT4-based evaluation prompts
@@
G.2.1 Prompt for evaluating helpfulness
@@
G.2.2 Prompt for evaluating harmlessness

### v2

@@ 7 Experiments
We empirically evaluate the effectiveness of SACPO in enhancing multiple criteria stepwise. This experiment focuses on improving helpfulness and safety (i.e., harmlessness). We use the same experimental setup as in Safe RLHF [15] wherever possible for fair comparisons. We employ the same SFT model (i.e., a reproduced version of Alpaca-7B [43]). This model is trained to function as a conversational assistant with proficient ability, which generates both benign and harmful responses. We utilize the PKU-SafeRLHF preference dataset [24] with more than 30,000 expert evaluations. Each record in this dataset presents a pair of responses to a specific prompt, and response pairs are ranked according to helpfulness and harmlessness. While the harmlessness of a response is determined by its neutrality concerning 14 different risk categories, the helpfulness is judged based on factors such as clarity, relevance, and overall quality.
@@
Implementations. In this experiment, we apply DPO and KTO for each alignment on helpfulness and safety (i.e., harmlessness). Specifically, we implement the following four variants of SACPO: DPO (H) $\rightarrow$ DPO (S), DPO (H) $\rightarrow$ KTO (S), KTO (H) $\rightarrow$ DPO (S), and DPO (S) $\rightarrow$ DPO (H), where H and S are abbreviations of helpfulness and safety (i.e., harmlessness). We use TRL [45] for implementing DPO and KTO. As for the parameter associated with the reverse KL divergence penalty, we first set $\beta$ and then test a wide range of values of $\beta/\lambda$. As a result, we set $\beta=0.1$ when helpfulness is the first alignment metric and $\beta=0.01$ otherwise. In addition, to evaluate the performance of P-SACPO presented in Section 6, we implement linear model merging [49] between the helpfulness-aligned model and conservatively safety-realigned model with $\beta/\lambda=0.01$ trained via DPO (H) $\rightarrow$ DPO (S). We use MergeKit [22] and test three different mixing ratios; that is, $q\in\{0.25,0.5,0.75\}$. For more implementation details (e.g., hyperparameters), see Appendix G.
Baselines. We evaluate the models trained via (P-)SACPO compared with the SFT model and those trained via Safe RLHF. Safe RLHF owns three models (i.e., beaver-7b-v1.0, -v2.0, and -v3.0), depending on the number of iterations regarding data collection and fine-tuning. Crucially, our (P-)SACPO optimizes LM policies under the same conditions as v1.0 and less favorable conditions than v2.0 and v3.0, in terms of the quality and quantity of data. For P-SACPO, the baseline method also includes naïve linear merging that simply averages the weights of DPO (H) and DPO (S).
@@
Evaluation. We use GPT-4 [1] to measure the helpfulness and harmlessness (i.e., safety) of the responses generated by the LM policies. We base our prompts on those in the Safe RLHF study with a slight adjustment in output format requirements to get more reliable evaluations (for more details, see Appendix G.4). Unlike the previous study, we employ two non-overlap sets of prompts to assess helpfulness and safety separately. We use prompts in the Safe RLHF study to evaluate safety, which has a high risk of inducing unsafe responses. For assessing helpfulness, we use prompts from the AlpacaEval dataset [28] that are unlikely to result in harmful content. Using two sets of prompts, we aim to ensure that our assessments of helpfulness and safety did not affect each other.
Results. Figure 2 shows the win rates of each model against the base SFT model.  [^2] Additional experimental results (e.g., Elo scores, statistical significance testing) are in Appendix G. First, Figure 2(a) illustrates the experimental results for DPO (H) $\rightarrow$ DPO (S), DPO (H) $\rightarrow$ KTO (S), and KTO (H) $\rightarrow$ DPO (S). We observe that DPO (H) and KTO (H) improve the performance on helpfulness at the first step. In DPO (H) $\rightarrow$ DPO (S) and KTO (H) $\rightarrow$ DPO (S), subsequent alignment for safety obtains a substantial improvement on harmlessness with a slight decrease in helpfulness. These models obtained by SACPO perform better than those obtained by Safe RLHF in terms of helpfulness and harmlessness. Notably, KTO (H) $\rightarrow$ DPO (S) performs well, which supports our main claim that different types of datasets or algorithms can be used for each alignment. Also, we observe that varying the $\beta/\lambda$ ratio allows us to fine-tune the equilibrium (i.e., a near Pareto-front) between helpfulness and safety. This result indicates the flexibility of the proposed method in obtaining a model with a desired trade-off between multiple criteria. However, DPO (H) $\rightarrow$ KTO (S) performs significantly worse than DPO (H) $\rightarrow$ DPO (S). We guess this is because KTO is inappropriate for the safety alignment, but we will leave it to future work to identify the detailed reasons. Figure 2(b) shows the effect of “order” of the stepwise alignment. This experimental result shows that we obtain the models with comparable performance regardless of the order of alignments, which is consistent with our theory (i.e., Remark 2). Finally, Figure 2(c) shows the effectiveness of P-SACPO proposed in Section 6, showing that P-SACPO performs better than the naïve method that simply averages the weights of DPO (H) and DPO (S). Linear model merging allows us to balance helpfulness and harmlessness by averaging reward-aligned and conservatively safety-realigned policies without optimizing new ones. Therefore, we can approximately find $\lambda^{\star}$ for the constrained LM policy optimization problem (10) with reduced computational time and stable learning.
@@
Figure 2: Win rate against the SFT model. H and S are abbreviations for helpfulness and safety (i.e., harmlessness), respectively. Crosses represent SFT and Safe RLHF, and blue circles represent models aligned with a single metric. (a) DPO (H) $\rightarrow$ DPO (S), DPO (H) $\rightarrow$ KTO (S), and KTO (H) $\rightarrow$ DPO (S). (b) DPO (S) $\rightarrow$ DPO (H). (c) P-SACPO based on linear model merging. Numbers represent $\beta/\lambda$ in (a) and (b) and $q$ in (c).
@@ 8 Conclusion
We have introduced SACPO, a simple algorithm for constrained language model policy optimization. SACPO takes a stepwise approach that sequentially aligns an LM policy using off-the-self alignment algorithms (e.g., DPO, KTO) and datasets for each metric. This procedure is theoretically justified and provides many practical benefits such as simplicity, stability, and flexibility. Our theoretical results include the upper bounds regarding near-optimality and safety constraint violations. Empirically, SACPO performs better than Safe RLHF in enhancing helpfulness and harmlessness, and we further show the effectiveness of a practical variant called P-SACPO based on linear model merging.
Limitations. SACPO has several limitations. First, while we evaluate SACPO using the models with 7B parameters, there is room for discussion on whether SACPO works well for state-of-the-art models with many more parameters. Second, although this paper focuses on safety alignment from the perspective of RLHF or DPO, it is more desirable to additionally incorporate the standard SFT as well as pre-check and post-check strategies. Finally, although SACPO is more efficient than existing methods, it still requires substantial computational cost or a large amount of high-quality data.
Broader Impacts. We believe SACPO contributes to the safety or trustworthiness of LLMs and will reduce the barrier to aligning future LLMs to enhance the benefits of AI while minimizing negative impacts. However, any LLMs are open to abuse, and models obtained by SACPO are not exceptions. Also, we must recognize that the core idea behind SACPO can be used to make LLMs more unsafe.
Appendix
@@ B.2 Lagrangian
B.3 Optimal Policy Can be Directly Obtained from Reward-aligned Policy
@@
Appendix C Proof of Lemma 1 and Lemma 2
Proof.
Proof.
Appendix D Proof of Theorem 1
@@
$=$ $\ \frac{Z_{r^{\star}}(x;\pi_{\mathrm{ref}})}{Z_{r^{\star}+\lambda^{\star}g^{\star}}(x;\pi_{\mathrm{ref}})}\underbrace{\frac{1}{Z_{r^{\star}}(x;\pi_{\mathrm{ref}})}\pi_{\mathrm{ref}}(y\mid x)\exp\left(\frac{1}{\beta}r^{\star}(x,y)\right)}_{=\pi^{\star}_{r^{\star}}(y\mid x)}\exp\left(\frac{\lambda^{\star}}{\beta}g^{\star}(x,y)\right)$
$=$ $\ \frac{1}{Y(x)}\pi^{\star}_{r^{\star}}(y\mid x)\exp\left(\frac{\lambda^{\star}}{\beta}g^{\star}(x,y)\right).$
@@ Appendix E Proof of Lemma 3
Recall Assumption 2. By Lemma 8 in Xiong et al. [50], the following inequality holds with probability at least $1-\delta$:
@@ F.1 Preliminary Lemmas
$\leq$ $\ c\cdot|\,g^{\star}(x,y)-\widehat{g}(x,y)\,|+|\,c-\widehat{\lambda}\,|\cdot|\,\widehat{g}(x,y)\,|.$
@@ Appendix G Details of the Experiments
G.1 Compute Resources
Our experiments were conducted in a workstation with Intel(R) Xeon(R) Silver 4316 CPUs@2.30GHz and 8 NVIDIA A100-SXM4-80GB GPUs.
G.2 Licenses
In the empirical experiment, we use the existing models or datasets. While we have properly cited the original papers in the main paper, we additionally list each license as follows.
  • Models
    • Alpaca-7B: CC By-NC-4.0
    • beaver-7b-v1.0, v-2.0, v-3.0: CC By-NC-4.0
  • Datasets
    • PKU-SafeRLHF: CC By-NC-4.0
    • Alpaca-Eval: CC By-NC-4.0
Our models are fine-tuned from Alpaca-7B using the PKU-SafeRLHF dataset; hence, we will make sure that the license of our models is also CC-By-NC-4.0 when we release them.
G.3 Hyper-parameters
@@
DPO (H) $\rightarrow$ DPO (S) $0.1$$\{0.01,0.025,0.05,0.1\}$
DPO (H) $\rightarrow$ KTO (safety) $0.1$$\{0.01,0.025,0.05,0.1\}$
KTO (H) $\rightarrow$ DPO (S) $0.1$$\{0.01,0.025,0.05,0.1\}$
DPO (S) $\rightarrow$ DPO (H) $0.01$$\{0.01,0.025,0.05,0.1\}$
G.4 GPT4-based evaluation prompts
@@
G.4.1 Prompt for evaluating helpfulness
@@
G.4.2 Prompt for evaluating harmlessness
@@
G.5 Elo scores
We provide additional experimental results. Specifically, we show the Elo scores of the models obtained by SACPO, compared to the SFT model and Safe RLHF. For a fair comparison, Safe RLHF corresponds to the beaver-7b-v1.0 model, which is trained using the same dataset as our models trained via SACPO and P-SACPO. The Elo scores are normalized so that the SFT model has a score of 1000.
First, Figure 3 show the Elo scores of DPO (H) $\rightarrow$ DPO (S) and DPO (H) $\rightarrow$ KTO (S). We observe that DPO (H) $\rightarrow$ DPO (S) performs better than the SFT and Safe RLHF. However, as with the win rate, the Elo scores of DPO (H) $\rightarrow$ KTO (S) are much worse than those of DPO (H) $\rightarrow$ DPO (S) or Safe RLHF.
Figure 3: Elo scores of DPO (H) $\rightarrow$ DPO (S) and DPO (H) $\rightarrow$ KTO (S).
Second, Figure 4 shows the Elo scores of (a) KTO (H) $\rightarrow$ DPO (S) and (b) DPO (S) $\rightarrow$ DPO (H). In both variants of SACPO, the Elo scores are better than the SFT model and Safe RLHF.
Figure 4: Elo scores of (a) KTO (H) $\rightarrow$ DPO (S) and (b) DPO (S) $\rightarrow$ DPO (H)
Finally, Figure 5 shows the Elo scores of the models trained by the P-SACPO based on the linear model merging. We see that P-SACPO based on the linear model merging performs better than SACPO with specific $\beta/\lambda$ in addition to the SFT model and Safe RLHF.
Figure 5: Elo scores of P-SACPO based on the linear model merging.
G.6 Significance Testing
We conduct statistical significance testing. We make GPT-4 evaluate each response three times. Table 3 shows the experimental results summarizing the mean and standard deviation ($1\sigma$) of the win rate against the SFT model. We observe that the standard deviation is fairly small. This indicates that our experimental results support the main claims of this paper in a statistically meaningful manner.
Table 3: Statistical significance testing of win rate against the SFT model. We compute the mean and standard deviation ($1\sigma$) across three random seeds.
Model Helpfulness ($\uparrow$) Harmlessness ($\uparrow$)
SFT $0.500\pm 0.000$ $0.500\pm 0.000$
Safe RLHF (beaver-7b-v1.0) $0.596\pm 0.013$ $0.692\pm 0.006$
DPO (H) with $\beta=0.1$ $0.792\pm 0.004$ $0.490\pm 0.009$
KTO (H) with $\beta=0.1$ $0.678\pm 0.009$ $0.458\pm 0.002$
DPO (S) with $\beta=0.01$ $0.605\pm 0.011$ $0.811\pm 0.006$
DPO (H) $\rightarrow$ DPO (S) with $\beta/\lambda=0.01$ $0.627\pm 0.011$ $0.793\pm 0.001$
DPO (H) $\rightarrow$ DPO (S) with $\beta/\lambda=0.025$ $0.685\pm 0.006$ $0.805\pm 0.001$
DPO (H) $\rightarrow$ DPO (S) with $\beta/\lambda=0.05$ $0.722\pm 0.005$ $0.697\pm 0.009$
DPO (H) $\rightarrow$ DPO (S) with $\beta/\lambda=0.1$ $0.731\pm 0.007$ $0.572\pm 0.005$
DPO (H) $\rightarrow$ KTO (S) with $\beta/\lambda=0.01$ $0.498\pm 0.011$ $0.454\pm 0.001$
DPO (H) $\rightarrow$ KTO (S) with $\beta/\lambda=0.025$ $0.479\pm 0.011$ $0.491\pm 0.004$
DPO (H) $\rightarrow$ KTO (S) with $\beta/\lambda=0.05$ $0.556\pm 0.009$ $0.530\pm 0.000$
DPO (H) $\rightarrow$ KTO (S) with $\beta/\lambda=0.1$ $0.660\pm 0.004$ $0.524\pm 0.007$
KTO (H) $\rightarrow$ DPO (S) with $\beta/\lambda=0.01$ $0.595\pm 0.013$ $0.856\pm 0.005$
KTO (H) $\rightarrow$ DPO (S) with $\beta/\lambda=0.025$ $0.725\pm 0.005$ $0.789\pm 0.002$
KTO (H) $\rightarrow$ DPO (S) with $\beta/\lambda=0.05$ $0.681\pm 0.011$ $0.730\pm 0.006$
KTO (H) $\rightarrow$ DPO (S) with $\beta/\lambda=0.1$ $0.734\pm 0.007$ $0.676\pm 0.012$
DPO (S) $\rightarrow$ DPO (H) with $\beta/\lambda=0.01$ $0.428\pm 0.011$ $0.696\pm 0.074$
DPO (S) $\rightarrow$ DPO (H) with $\beta/\lambda=0.025$ $0.638\pm 0.004$ $0.761\pm 0.001$
DPO (S) $\rightarrow$ DPO (H) with $\beta/\lambda=0.05$ $0.672\pm 0.008$ $0.738\pm 0.011$
DPO (S) $\rightarrow$ DPO (H) with $\beta/\lambda=0.05$ $0.672\pm 0.006$ $0.807\pm 0.007$
P-SACPO with $q=0.25$ $0.798\pm 0.015$ $0.545\pm 0.002$
P-SACPO with $q=0.5$ $0.787\pm 0.005$ $0.651\pm 0.004$
P-SACPO with $q=0.75$ $0.764\pm 0.005$ $0.768\pm 0.007$
naïve model merging with $q=0.75$ $0.816\pm 0.007$ $0.588\pm 0.008$
naïve model merging with $q=0.5$ $0.759\pm 0.006$ $0.617\pm 0.004$
naïve model merging with $q=0.5$ $0.694\pm 0.009$ $0.690\pm 0.003$